OpenStudy (anonymous):

Use elimination to solve the system of equations. Find the ordered pairs. { 2x-5=-5 -2x+8y= -58

OpenStudy (hoblos):

are you missing a (y)

hero (hero):

Should it be: 2x-5y=-5 -2x+8y= -58

hero (hero):

If so, then: 3y = -63 y = -21

OpenStudy (hoblos):

2x-5y=-5 -2x+8y= -58 add 3y=-63 y=-21 x=-55